OpenStudy (anonymous):

Find the factors of w2 − 16.

OpenStudy (anonymous):

= (w+4)(w-4)

OpenStudy (anonymous):

ohhh i see thanks :D

OpenStudy (anonymous):

Use: \[a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

Welcome dear..